OpenStudy (anonymous):

Discuss the symmetry between the laws of Charles Coulomb and Sir Isaac Newton regarding forces Matt forgot to put the fabric softener in the wash. As his socks tumbled in the dryer, they became charged. If a small piece of lint with a charge of +1.25 E -19 C is attracted to the socks by a force of 3.0 E -9 N, what is the magnitude of the electric field at this location? Suppose an electric field with a magnitude of 2.8 E 4 N/C at a certain location, points due south. If a charge of -4.0 E -6 C is brought to the location, what is the electrostatic force that acts on this c

OpenStudy (anonymous):

E = F/q E = the electric field F = the electrostatic force experienced by a charge q = the magnitude of the electrical charge If the electrostatic force is expressed in newtons (N), and the charge magnitude is expressed in coulombs (C), then the SI unit for the magnitude of the electric field is newtons/coulomb (N/C). The electric field is a vector quantity requiring a direction for a complete description. The direction is the same direction as the force on a positive test charge.

OpenStudy (anonymous):

Coulomb's law has the force on one charged body due to another equal and opposite to the force on that other charged body due to the first. Force, F = (charge ) (field) = q E, so get E from F/q. Again, use F = q E, now getting force. F in newtons, q in coulombs, E in volts/m = N/C.

OpenStudy (anonymous):

how does this compare to newtons forces

OpenStudy (anonymous):

Newton's law of gravitation has force proportional to the product of the masses of the two bodies and inversely proportional to the square of the distance between them. Coulomb's law is similar, with the charges replacing the masses; however, you can have repulsion (+)(+) or (-)(-) with electrostatics but only attraction with gravitation.